An air standard dual-combustion cycle has a mean effective pressure of 10 bar. The minimum pressure and temperature are 1 bar and 17 ?C, respectively, and the compression ratio is 16/1. Calculate the maximum cycle temperature when the cycle efficiency is 60%. The maximum cycle pressure is 60 bar.
1
Expert's answer
2021-09-12T00:30:26-0400

"P_1=1bar"

"V_1\/V_2=16"

"P_3=60 bar"

"C_v=0.718"

"C_p=1.005"

"T_2=T_1\times(V_1/V_2)^{r-1}"

"T_2=879K"

"P_2=P_1\times(V_1/V_2)^r"

"P_2=48.5 bar"

"T_2=T_2 \times P_3/P_2"

"=879\times60/48.5"

"T_3=1087K"

Now,

"V_4\/V_3=T_4\/T_3"

"V_4/V_3\times T_3 =T_4"

"T_4=1087\times (V_4/V_3)" ——————(*)

"T_5=0.3298\times T_4 \times (V_4/V_3)^{1.4-1}"

"T_5=0.020126\times T_41.4"

"=1-\frac{C_v(T_5-T_2)}{C_v(T_5-T_2)+C_p(T_4-T_3)}"

by the equation putting values;

"T_4=1959k"

when the cycle efficiency is 60%